OpenStudy (winsteria):

FAN AND MEDAL Use slope formula,, to find the slope of a line that passes through the points (–3, 8) and (4, –6). m = ? 1)-7 2)-2 3)1/2 4)2 Use slope-intercept form, y = mx + b, to find the y-intercept (b) of the line. b = ? 1)-14 2)-8 3)2 4)7 What is the new equation written in slope-intercept form, y = mx + b? 1)y=(1/2)x-8 2)y=-2x+2 3)y=2x-14 4)y=-2x-8

pooja195 (pooja195):

1) Use the slope formula \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

Mehek (mehek14):

\(\dfrac{-6-8}{4-(-3)}\)

OpenStudy (winsteria):

-14 _____ ? 7

Mehek (mehek14):

which is simplified to?

Mehek (mehek14):

divide -14/7 = ?

OpenStudy (winsteria):

-2

Mehek (mehek14):

correct so slope = -2

Mehek (mehek14):

now we need the y-intercept so far, the equation is y = -2x + b b = y-intercept we will use the ordered pair (-3,8) x = -3 y = 8 8 = -2 * -3 + b 8 = 6 + b 8 - 6 = 2 2 = b

OpenStudy (winsteria):

I get that thank you

OpenStudy (winsteria):

what do I do for the last question though?

Mehek (mehek14):

if you have slope -2 and y-intercept 2 what do you think goes best?

OpenStudy (winsteria):

y=-2x+2

Mehek (mehek14):

yup

OpenStudy (winsteria):

Thank you

Mehek (mehek14):

yw :)